interest = 3%Q2. The Department of Engineering is contemplating the purchase of a top-of-the- line PCB drilling machine to be used in its
Answer #1

Given,

Interest rate = 3%

Price of the machine = $5,000

Market value at the end of life (10 years) = $S

Depreciation rate per year under SL depreciation = (Price - Market value at the end of life)/Useful life = ($5,000 - $S)/10 = 500 - 0.1S

Total depreciation up to the EOY 8 = 8 * Depreciation rate per year = 8 * (500 - 0.1S) = 4000 - 0.8S

Market value at the EOY 8 = Price - Total depreciation up to the EOY 8 = 5000 - (4000-0.8S) = 1000 + 0.8S

Annual fee paid by the students = $500

From the compound interest table, we obtain

(P/A, 3%, 8) = 7.020

(P/F, 3%, 8) = 0.7894

Present worth of annual fee = $500*(P/A, 3%, 8) = $500 * 7.020 = $3510

Present worth of the machine expenses = -$5000 + (1000+0.8S)*(F/P, 3%, 8) = -$5000 + (1000+0.8S)*0.7894 = -5000 + 789.4 + 0.6315S = -4210.6 + 06315S

In order to break-even, the net present worth of the machine = 0

=> PW of annual fee + PW of the machine expenses = 0

=> 3510 - 4210.6 + 0.6315S = 0

=> 0.6315S = 700.6

=> S = 1109.4

Ans: S = 1109.4